Three resistors 2ohm, 4ohm, and 5ohm are combined in parallel. What is the total resistance of the combination? 

Grade:12

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

( 0.5+0.25+0.2)^(-1)

= 100/95= 20/19 ohms

Regards
Vikas TU
4 years ago
Dear student 
Please refer the link below 
God Luck 
 Chheers